hi! what is the vertices of the ellipse equation: (x-2)^2+((y+1)^2)/9=1
Answer by lwsshak3(11628)

What is the vertices of the ellipse equation: (x-2)^2+((y+1)^2)/9=1
**
(x-2)^2+((y+1)^2)/9=1
Standard form of an equation for ellipse with vertical major axis: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
For given equation:
center: (2,-1)
a^2=9
a=3
vertices: (2,-1±a)=(2,-1±3)=(2,-4) and (2,2)